# Detectability of dark matter subhalo impacts in Milky Way stellar streams

### v2

@@ Abstract
Stellar streams are a promising way to probe the gravitational effects of low-mass dark matter (DM) subhalos. In recent years, there has been a remarkable explosion in the number of stellar streams detected in the Milky Way, and hundreds more may be discovered with future surveys such as LSST. Studies of DM subhalo impacts on streams have so far focused on a few of the thinnest and brightest streams, and it is not known how much information can be gained from the others. In this work, we develop a method to quickly estimate the minimum detectable DM subhalo mass of a given stream, where subhalo mass here refers to the total mass of a Plummer sphere. Our work is based on an analytic model for subhalo impacts on circular streams, which allows us to model streams with a wide range of properties including width, length, distance, and stellar density. We consider several observational scenarios, based on current and future surveys including Gaia, DESI, Via, and LSST. We find that at 95% confidence level, a stream like GD-1 has a minimum detectable subhalo mass of $\sim 6\times 10^{6}~{}\mathrm{M}_{\odot}$ in Gaia data and $\sim 8\times 10^{5}~{}\mathrm{M}_{\odot}$ with LSST 10 year sensitivity. Applying our results to confirmed Milky Way streams, we rank order them by their sensitivity to DM subhalos and identify promising ones for further study.
@@ 1 Introduction
The predictions of the standard cosmological model, $\Lambda$CDM, match well with measurements of the matter power spectrum $P(k)$ at large scales ($k\lesssim 10~{}\mathrm{h/Mpc}$). At small scales, however, we currently lack reliable measurements. Various dark matter (DM) models with different properties can have drastically different behaviors at small scales ($k\gtrsim 10~{}\mathrm{h/Mpc}$(Bechtol et al., 2023). Probing small scales will test the paradigm of cold dark matter in a new regime, and could reveal hints for the underlying particle physics of dark matter.
@@
In the late universe, density perturbations with $k\gtrsim 10~{}\mathrm{h/Mpc}$ have become nonlinear and clustered into bound Galactic-scale structures and substructures. Measuring DM subhalo abundances in the range $10^{5}\,\mathrm{M}_{\odot}-10^{8}\,\mathrm{M}_{\odot}$ can effectively put constraints on small scales ($10^{1}-10^{3}~{}\mathrm{h/Mpc}$) of the matter power spectrum. Several observational probes have been proposed to determine the abundance of DM subhalos in this mass range, including ultra-faint dwarf galaxies (e.g. Nadler et al. (2024)), strong gravitational lensing (e.g. Keeley et al. (2024)) and stellar streams (e.g. Banik et al. (2021a)).
@@
The focus of this paper is on the potential for stellar streams to probe these low mass DM subhalos. Stellar streams are formed from the tidal debris of dwarf galaxies and globular clusters in the Galactic halo (Lynden-Bell & Lynden-Bell, 1995; Bonaca & Price-Whelan, 2025). The resulting structures, especially those formed from globular clusters, are thin and dynamically cold, making them sensitive to the gravitational effects of low-mass DM subhalos (Ibata et al., 2002; Johnston et al., 2002). A subhalo passing near a stream, referred to as an impact, induces a perturbation, with measurements of streams potentially providing a record of past impacts.
@@
The most commonly discussed signature of a subhalo impact is a density fluctuation or gap (Yoon et al., 2011; Carlberg, 2012). The rates for impacts to cause gaps in streams have been studied both analytically and in simulations (Carlberg, 2012; Barry et al., 2023; Menker & Benson, 2024; Adams et al., 2024), and it is estimated that a gap from a subhalo as low as a few ×$\times 10^{5}\mathrm{M}_{\odot}$ could be detectable with LSST photometry (Drlica-Wagner et al., 2019). Rather than identifying individual gaps, the cumulative history of subhalo impacts can also be probed with the power spectrum of density fluctuations (Bovy et al., 2017). This technique similarly could be sensitive to subhalos down to $10^{5}\mathrm{M}_{\odot}$. Some of the above density features have already been found in well-studied streams such as GD-1, and shown to be consistent with impacts from DM subhalos at the $\sim 10^{7}\mathrm{M}_{\odot}$ mass scale (Bonaca et al., 2019a; Banik et al., 2021b, a). However, other effects also give rise to perturbations in streams, including epicyclic density variations (Ibata et al., 2020), progenitor disruption (Webb & Bovy, 2019), and other satellites and baryonic structures in the Milky Way (de Boer et al., 2020; Bonaca et al., 2020).
@@
In addition to gaps, DM subhalos leave other signatures such as velocity perturbations and angular deflections. Accounting for these can add valuable information to help disentangle the potential sources of a stream perturbation. Previous works (Erkal & Belokurov, 2015b; Hilmi et al., 2024) have developed frameworks to infer impact properties from observed density and kinematic data. Precise kinematics would aid in reconstruction of a perturber’s orbit and mass, in principle allowing for its identification. Photometric surveys like Gaia (Gaia Collaboration et al., 2021) and LSST (LSST Science Collaboration et al., 2009) can provide proper motions for the stars. Meanwhile, spectroscopic surveys like S5 (Li et al., 2022), DESI (Cooper et al., 2023), Via [^1] https://via-project.org, 4MOST (de Jong et al., 2012), and WEAVE (Dalton et al., 2012) can provide confirmation of stream stars and line-of-sight velocities.
@@
A limitation of prior studies is that they have all focused on DM subhalo detection in a few exemplar streams, in particular assuming a cold and long stream, of which the prototypical example is GD-1 (Grillmair & Dionatos, 2006). This ideal stream is certainly well-suited to the problem. At the same time, there has been a rapid growth in the number of Milky Way stellar streams identified in the era of Gaia (Bonaca & Price-Whelan, 2025) with over 100 confirmed (Ibata et al., 2024). There may be many more in Gaia data waiting to be confirmed (Shih et al., 2024) and hundreds more discovered with LSST (Pearson et al., 2024). It is essential to understand how much information we can gain from all of the streams combined, and which of the observed streams are the most promising.
@@
In this paper, we build a statistical framework for rapidly estimating the detectability of DM subhalo impacts in Milky Way streams, accounting for stream properties such as density, internal dispersion or width, and distance. Our focus is on the minimum detectable DM subhalo mass, as this is the quantity most closely tied to the matter power spectrum. The DM subhalo is modeled here as a Plummer sphere, with the mass defined as the total mass. We consider the detectability of a single strong impact, motivated by previous studies estimating that there are $O(1)$ impacts leading to a detectable gap over the lifetime of the stream (Erkal et al., 2016; Menker & Benson, 2024; Adams et al., 2024). We systematically explore how stream properties affect the minimum detectable subhalo mass in this impact, and use our resulting model to rank order the currently-known stellar streams of the Milky Way. Our study therefore paves the way for considering the population of detectable impacts in Milky Way stellar streams.
@@
Streams in the Milky Way can be characterized by a number of properties, and are each on a unique orbit. Rather than modeling all streams in detail, we will consider an idealized model of a circular stream with the observer at the Galactic Center. The streams are characterized by intrinsic properties, in particular stream width in angle $\sigma_{\theta}$, total number of stars per unit angle $\lambda$, distance to the Galactic Center $r_{0}$ and length in angle $l$. When the subhalo is modeled as a Plummer potential, there is a simple analytic approximation for the effect of a subhalo impact on the circular stream (Erkal & Belokurov, 2015a). The advantage of this approach is that it enables rapid simulation of impacted streams, evaluation of likelihoods, and extensive exploration over subhalo and stream properties. It also provides a clean way to illustrate the basic dependencies.
The detectability of subhalos depends of course on the available observations. We will evaluate four observational scenarios, ranging from idealized Gaia-era data all the way to future LSST data combined with the most precise spectroscopic surveys. With this, we build up a general analytic formula for minimum detectable subhalo mass as a function of stream properties, and for different observational scenarios. We apply this model to the known catalog of stellar streams and provide a ranking of the most promising candidates.
@@
In Sec. 2, we describe the setup for modeling subhalo impacts on a circular stream, which relies on a combination of the analytic model and simulated streams to obtain stream dispersion. In Sec. 3, we discuss how observational errors and limits are accounted for, introducing an effective per-star error that folds in all effects. We identify four benchmark observational scenarios. The statistical framework is presented in Sec. 4, along with the results for minimum detectable subhalo mass. The expected precision on subhalo mass is also given. Our main result for detectability as a function of stream properties is given in Sec. 4.2. In Sec. 5, we apply our result to streams of the Milky Way. A discussion of how our results would change in more realistic stream models can be found in Sec. 6, and conclusions in Sec. 7.
@@ 2.1 Analytic model for subhalo impact
Figure 1: Left: The geometry of a subhalo passing by a stellar stream. The stream is moving in a circular orbit in the $xy$ plane of the galaxy. At the point of closest approach, the subhalo passes by the stream with distance $b$, and velocity $(w_{r},w_{t},w_{z})$ in the Galactic frame. Right: A comparison between the observables and errors generated from simulated perturbed streams (blue) versus using the analytic model combined with errors from simulated unperturbed streams (orange). The blue dots show the simulated stars in the perturbed stream. The stream is generated with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$ and $\lambda=10~{}\mathrm{deg}^{-1}$. The subhalo impact is parametrized by $M_{\mathrm{sh}}=10^{7.5}~{}\mathrm{M}_{\odot}$, $r_{s}=0.91~{}\mathrm{kpc}$, $b=0~{}\mathrm{kpc}$, $t=315~{}\mathrm{Myr}$, $w_{r}=0~{}\mathrm{km/s}$, $w_{t}=0~{}\mathrm{km/s}$ and $w_{z}=180~{}\mathrm{km/s}$.
We make use of the model developed in Erkal & Belokurov (2015a, b) to approximate the subhalo impact on stellar streams. In this model, all stream stars are assumed to initially be on the progenitor’s orbit, and at some point in time, they instantaneously receive a velocity kick due to the gravitational effect of the passing DM subhalo. Treating these kicks as small perturbations, analytic results for the perturbed orbits of the stream stars can be obtained. This gives a model for the perturbed stream as a function of time after the impact.
@@
During the subhalo flyby, the stream is approximated to be moving with a constant velocity along a straight line. We parameterize location along the stream arm at this time by an angle $\psi_{0}$, where $\psi_{0}=0$ is the point of closest approach. Treating the impact with the impulse approximation, then stars along the stream get velocity kicks given by (Erkal & Belokurov, 2015b)
@@
From these velocity kicks, Erkal & Belokurov (2015b) derives analytic formulae for 6 observables today as functions of $\psi_{0}$:
@@
Here $\psi_{0}$ is the angle along the stream at the time of impact, while we are interested in these observables as functions of the observed angle along the stream today, which is denoted by $\psi$. We transform the functions of $\psi_{0}$ above to functions of $\psi$ using the formula for $\psi(\psi_{0},t)$ in Erkal & Belokurov (2015b).
@@ 2.2 Simulated streams and stream characteristics
In the analytic model, the stream is on a single orbit, which therefore assumes an infinitely thin stream. Now considering simulated streams with finite dispersion, the analytic model gives the expected value of observables along the stream. We also need to know the errors for the 6 observables. In this section, we deal with the statistical errors, which are controlled by the internal dispersion of the stream and number density along the stream. For this, we generate simulated streams and calculate their internal dispersion.
In order to introduce errors for observables, we split the stream into evenly sized angular bins in $\psi$. We choose to use 2 degree bins throughout the paper. This ensures we have as many stars per bin as possible while not losing the perturbation features across bins. To keep the notation concise, we will use $X$ to denote the shift in any of the position or velocity quantities $\Delta r$, $\Delta\theta$, $\Delta v_{r}$, $\Delta v_{t}$ and $\Delta v_{z}$. For a given angular bin along $\psi$ with $n$ stars in the bin, the observable is $\overline{X}=\sum_{i}X_{i}/n$ with $X_{i}$ being the individual measurements. All observed quantities are defined with respect to the unperturbed stream (e.g. $\Delta r_{i}=r_{i}-r_{0}$). The error on the mean value of $X$ is given by
@@
where $\sigma_{X}$ is the error or dispersion for individual stars, which can be obtained from the simulated streams. Note that we haven’t included any observational errors here and will introduce them in more detail in Sec. 3. In addition, to generalize these observables to arbitrary streams, one also has to account for errors on the baseline unperturbed model achieved from fitting the stream orbit; we assume that such errors are small and do not include them.
Finally, for the observable density ratio $\rho/\rho_{0}$, the Poissonian error is simply
@@
Again, $\rho_{0}$ is the density from the unperturbed analytic model, which is a constant across bins and does not contribute to the error.
@@
With the above results, we can generate mock data by sampling the observables in each $\psi$ bin. The mean is given by the analytic model and the errors are given by Eq. 5 and 6. Observational errors can also be added to this procedure, and we will consider them in detail in Sec. 3. Note that this process assumes that the dispersion along the stream is not substantially modified by the presence of the subhalo, which we will validate later in Sec. 2.3.
Figure 2: Dependence of some key observables, $\overline{\Delta\theta}$ and $\overline{\Delta v_{r}}$, on stream properties. The subhalo impact is the same as the right panel of Fig. 1. In each column, we vary one stream property and fix the others to demonstrate how a given property affects either the size of the error bars or the shape of the signal. The default stream properties are $\sigma_{\theta}=0.2$ deg, $r_{0}=10~{}\mathrm{kpc}$ and $\lambda=32~{}\mathrm{deg}^{-1}$. Signals here are only calculated from the analytic model and do not account for the errors. Error bars here are only due to stream dispersion. When observational errors are included (Sec. 3), changing $r_{0}$ will also affect the error bars.
@@
and the dispersion in the radial, tangential and $z$ velocities are proportional to the ratio of tidal radius to the orbital radius,
$\sigma_{v_{r}}\propto\sigma_{v_{t}}\propto\sigma_{v_{z}}\propto\frac{r_{\mathrm{tidal}}}{r_{0}},$ (8)
where the tidal radius is
@@
$\sigma_{r}\propto\left(\frac{M_{\mathrm{prog}}}{r_{0}}\right)^{1/3}r_{0}$ (10)
and
@@
From Eq. 5 and Eq. 6, the error also depends on the number of stars per bin $n$, or essentially the number of stars per unit angle $\lambda$. Following Fardal et al. (2015), the stream density in the central part of the stream is controlled by the total number of stream stars $N_{\mathrm{stars}}$ as,
@@
In addition to errors, the detectability also depends on the signals (observables) of a subhalo impact. From Eqs. 1-3, we see that the observables depend on the stream’s distance to the Galactic Center $r_{0}$. The length of the stream is also important for the full signal to be present, since the impact is most relevant over a region of length $\sim w_{\mathrm{rel}}\sqrt{b^{2}+r_{s}^{2}}/w_{\perp}$. From Fardal et al. (2015), the angular length of the stream $l$ is controlled by the age of the stream $T_{\mathrm{form}}$ and the velocity dispersion in the tangential direction $\sigma_{v_{t}}$:
@@
 $r_{0}$ distance to the Galactic Center $5-40$ kpc

### v3

@@ Abstract
Stellar streams are a promising way to probe the gravitational effects of low-mass dark matter (DM) subhalos. In recent years, there has been a remarkable explosion in the number of stellar streams detected in the Milky Way, and hundreds more may be discovered with future surveys such as LSST. Studies of DM subhalo impacts on streams have so far focused on a few of the thinnest and brightest streams, and it is not known how much information can be gained from the others. In this work, we develop a method to quickly estimate the minimum detectable DM subhalo mass of a given stream, depending on its width, length, distance, and stellar density. We use an analytic model for the impacts and apply a test statistic to determine whether they are detectable. We consider several observational scenarios, based on current and future surveys including Gaia, DESI, Via, and LSST. We find that at 95% confidence level, a stream like GD-1 has a minimum detectable subhalo mass of $\sim 6\times 10^{6}~\mathrm{M}_{\odot}$ in Gaia data and $\sim 8\times 10^{5}~\mathrm{M}_{\odot}$ with LSST 10 year sensitivity. Applying our results to confirmed Milky Way streams, we rank order them by their sensitivity to DM subhalos and identify promising ones for further study.
@@ 1 Introduction
[…]
@@
In the late universe, density perturbations with $k\gtrsim 10~\mathrm{h/Mpc}$ have become nonlinear and clustered into bound Galactic-scale structures and substructures. Measuring DM subhalo abundances in the range $10^{5}\,\mathrm{M}_{\odot}-10^{8}\,\mathrm{M}_{\odot}$ can effectively put constraints on small scales ($10^{1}-10^{3}~\mathrm{h/Mpc}$) of the matter power spectrum (Bullock and Boylan-Kolchin, 2017). Several observational probes have been proposed to determine the abundance of DM subhalos in this mass range, including ultra-faint dwarf galaxies (e.g. Macciò and Fontanot (2010); Nadler et al. (2024)), strong gravitational lensing (e.g. Gilman et al. (2020); Keeley et al. (2024)) and stellar streams (e.g. Ibata et al. (2002); Johnston et al. (2002); Banik et al. (2021b)).
@@
The focus of this paper is on the potential for stellar streams to probe these low mass DM subhalos. Stellar streams are formed from the tidal debris of dwarf galaxies and globular clusters in the Galactic halo (Lynden-Bell and Lynden-Bell, 1995; Johnston et al., 1996; Helmi and White, 1999; Bonaca and Price-Whelan, 2025). The resulting structures, especially those formed from globular clusters, are thin and dynamically cold, making them sensitive to the gravitational effects of low-mass DM subhalos (Ibata et al., 2002; Johnston et al., 2002). A subhalo passing near a stream, referred to as an impact, induces a perturbation, with measurements of streams potentially providing a record of past impacts.
@@
The most commonly discussed signature of a subhalo impact is a density fluctuation or gap (Yoon et al., 2011; Carlberg, 2012). The rates for impacts to cause gaps in streams have been studied both analytically and in simulations (Carlberg, 2012; Erkal et al., 2016; Barry et al., 2023; Menker and Benson, 2024; Adams et al., 2024), and it is estimated that a gap from a subhalo as low as a few $\times 10^{5}\mathrm{M}_{\odot}$ could be detectable with LSST photometry (Drlica-Wagner et al., 2019). Rather than identifying individual gaps, the cumulative history of subhalo impacts can also be probed with the power spectrum of density fluctuations (Bovy et al., 2017). This technique similarly could be sensitive to subhalos down to $10^{5}\mathrm{M}_{\odot}$. Some of the above density features have already been found in well-studied streams such as GD-1, and shown to be consistent with impacts from DM subhalos at the $\sim 10^{7}\mathrm{M}_{\odot}$ mass scale (Bonaca et al., 2019b; Banik et al., 2021a, b). However, other effects also give rise to perturbations in streams, including epicyclic density variations (Küpper et al., 2012; Ibata et al., 2020), progenitor disruption (Webb and Bovy, 2019), and other satellites and baryonic structures in the Milky Way (Erkal et al., 2017; Pearson et al., 2017; Banik and Bovy, 2019; de Boer et al., 2020; Bonaca et al., 2020).
@@
In addition to gaps, DM subhalos leave other signatures such as velocity perturbations and angular deflections. Accounting for these can add valuable information to help disentangle the potential sources of a stream perturbation. Previous works (Erkal and Belokurov, 2015b; Hilmi et al., 2024) have developed frameworks to infer impact properties from observed density and kinematic data. Precise kinematics would aid in reconstruction of a perturber’s orbit and mass, in principle allowing for its identification. Photometric surveys like Gaia (Gaia Collaboration et al., 2021) and LSST (LSST Science Collaboration et al., 2009) can provide proper motions for the stars. Meanwhile, spectroscopic surveys like S5 (Li et al., 2022), DESI (Cooper et al., 2023), Via [^1] https://via-project.org, 4MOST (de Jong et al., 2012), and WEAVE (Dalton et al., 2012) can provide confirmation of stream stars and line-of-sight velocities.
@@
A limitation of prior studies is that they have all focused on DM subhalo detection in a few exemplar streams, in particular assuming a cold and long stream, of which the prototypical example is GD-1 (Grillmair and Dionatos, 2006). This ideal stream is certainly well-suited to the problem. At the same time, there has been a rapid growth in the number of Milky Way stellar streams identified in the era of Gaia (Bonaca and Price-Whelan, 2025) with over 100 confirmed (Mateu, 2023; Ibata et al., 2024). There may be many more in Gaia data waiting to be confirmed (Shih et al., 2024) and hundreds more discovered with LSST (Pearson et al., 2024). It is essential to understand how much information we can gain from all of the streams combined, and which of the observed streams are the most promising.
@@
In this paper, we build a statistical framework for rapidly estimating the detectability of DM subhalo impacts in Milky Way streams, accounting for stream properties such as density, internal dispersion or width, and distance. Our focus is on the minimum detectable DM subhalo mass, as this is the quantity most closely tied to the matter power spectrum. The subhalo mass refers to the total mass assuming that the subhalo is modeled as a Plummer sphere. We consider the detectability of a single strong impact, motivated by previous studies estimating that there are $O(1)$ impacts leading to a detectable gap over the lifetime of the stream (Erkal et al., 2016; Menker and Benson, 2024; Adams et al., 2024). We systematically explore how stream properties affect the minimum detectable subhalo mass in this impact, and apply our resulting model to currently-known stellar streams of the Milky Way.
@@
Streams in the Milky Way can be characterized by a number of properties, and are each on a unique orbit. Rather than modeling all streams in detail, we will consider an idealized model of a circular stream with the observer at the Galactic Center. The streams are characterized by intrinsic properties, in particular stream width in angle $\sigma_{\theta}$, total number of stars per unit angle $\lambda$, distance to the Galactic Center $r_{0}$ and length in angle $l$. When the subhalo is modeled as a Plummer potential, there is a simple analytic approximation for the effect of a subhalo impact on the circular stream (Erkal and Belokurov, 2015a). The advantage of this approach is that it enables rapid simulation of impacted streams, evaluation of likelihoods, and extensive exploration over subhalo and stream properties. It also provides a clean way to illustrate the basic dependencies.
The detectability of subhalos depends of course on the available observations. We will evaluate four observational scenarios: a) idealized Gaia-era data, b) Gaia data combined with spectroscopic data from DESI, c) LSST single epoch sensitivity with spectroscopic data from Via and d) LSST 10 year sensitivity with spectroscopic data from Via. The LSST 10 year sensitivity data is expected to include 10 times more stars at a distance of 10 kpc compared to Gaia (see Fig. 3). Via will improve the accuracy of radial velocity by a factor of 10 compared to DESI (see top left panel of Fig. 4).
With this, we build up a general analytic formula for minimum detectable subhalo mass as a function of stream properties, and for different observational scenarios. We apply this model to the known catalog of stellar streams and provide a ranking of the most promising candidates. We emphasize this ranking is not the complete picture for the most promising streams, since detection prospects also depend on the expected rate of subhalo impacts as well as baryonic backgrounds. However, our study provides a foundation for considering the population of detectable impacts in Milky Way stellar streams.
@@
In Sec. 2, we describe the setup for modeling subhalo impacts on a circular stream, which relies on a combination of the analytic model and simulated streams to obtain stream dispersion. In Sec. 3, we discuss how observational errors and limits are accounted for, introducing a survey-effective error that folds in all effects. We identify four benchmark observational scenarios. The statistical framework is presented in Sec. 4, along with the results for minimum detectable subhalo mass. The expected precision on subhalo mass is also given. Our main result for detectability as a function of stream properties is given in Sec. 4.2. In Sec. 5, we apply our result to streams of the Milky Way. A discussion of how our results would change in more realistic stream models can be found in Sec. 6, and conclusions in Sec. 7.
@@ 2.1 Analytic model for subhalo impact
Figure 1: Left: The geometry of a subhalo passing by a stellar stream. The stream is moving in a circular orbit in the $xy$ plane of the galaxy. At every point on the stream, the stars are moving along the tangential $\hat{t}$ direction. At the point of closest approach, the subhalo passes by the stream with distance $b$, and velocity $(w_{r},w_{t},w_{z})$ in the Galactic frame. Right: A comparison between the observables and uncertainties from simulated perturbed streams (blue) versus using the analytic model combined with random Gaussian noises generated based on uncertainties of simulated unperturbed streams (orange). The blue dots show the simulated stars in the perturbed stream. The stream is generated with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$ and $\lambda=10~\mathrm{deg}^{-1}$. The subhalo impact is parametrized by $M_{\mathrm{sh}}=10^{7.5}~\mathrm{M}_{\odot}$, $r_{s}=0.91~\mathrm{kpc}$, $b=0~\mathrm{kpc}$, $t=315~\mathrm{Myr}$, $w_{r}=0~\mathrm{km/s}$, $w_{t}=0~\mathrm{km/s}$ and $w_{z}=180~\mathrm{km/s}$.
We make use of the model developed in Erkal and Belokurov (2015a, b) to approximate the subhalo impact on stellar streams. In this model, all stream stars are assumed to initially be on the progenitor’s orbit, and at some point in time, they instantaneously receive a velocity kick due to the gravitational effect of the passing DM subhalo. Treating these kicks as small perturbations, analytic results for the perturbed orbits of the stream stars can be obtained. This gives a model for the perturbed stream as a function of time after the impact.
@@
During the subhalo flyby, the stream is approximated to be moving with a constant velocity along a straight line. We parameterize location along the stream arm at this time by an angle $\psi_{0}$, where $\psi_{0}=0$ is the point of closest approach. Treating the impact with the impulse approximation, then stars along the stream get velocity kicks given by (Erkal and Belokurov, 2015b)
@@
From these velocity kicks, Erkal and Belokurov (2015b) derives analytic formulae for 6 observables today as functions of $\psi_{0}$:
@@
Here $\psi_{0}$ is the angle along the stream at the time of impact, while we are interested in these observables as functions of the observed angle along the stream today, which is denoted by $\psi$. We transform the functions of $\psi_{0}$ above to functions of $\psi$ using the formula for $\psi(\psi_{0},t)$ in Erkal and Belokurov (2015b).
@@ 2.2 Simulated streams and stream characteristics
In the analytic model, the stream is on a single orbit, which therefore assumes an infinitely thin stream. For streams with a non-zero dispersion, we use this analytic model to predict the expected mean value of the observables along the stream. We also need to know the uncertainties for the 6 observables. In this section, we deal with the statistical uncertainty, which is controlled by the internal dispersion of the stream and number density along the stream. For this, we generate simulated streams and calculate their internal dispersion.
In order to introduce uncertainties for observables, we split the stream into evenly sized angular bins in $\psi$. We choose to use 2 degree bins throughout the paper. This ensures we have as many stars per bin as possible while not losing the perturbation features across bins. To keep the notation concise, we will use $X$ to denote the shift in any of the position or velocity quantities $\Delta r$, $\Delta\theta$, $\Delta v_{r}$, $\Delta v_{t}$ and $\Delta v_{z}$. For a given angular bin along $\psi$ with $n$ stars in the bin, the observable is $\overline{X}=\sum_{i}X_{i}/n$ with $X_{i}$ being the individual measurements. All observed quantities are defined with respect to the unperturbed stream (e.g. $\Delta r_{i}=r_{i}-r_{0}$). The uncertainty on the mean value of $X$ is given by
@@
where $\sigma_{X}$ is the dispersion of stars in the stream, which can be obtained from the simulated streams. Note that we have not included any observational errors here and will introduce them in more detail in Sec. 3. In addition, to generalize these observables to arbitrary streams, one also has to account for uncertainties on the baseline unperturbed model achieved from fitting the stream orbit; we assume that such uncertainties are small and do not include them.
Finally, for the observable density ratio $\rho/\rho_{0}$, the Poissonian uncertainty is simply
@@
Again, $\rho_{0}$ is the density from the unperturbed analytic model, which is a constant across bins and does not contribute to the uncertainty.
@@
With the above results, we can generate mock data by sampling the observables in each $\psi$ bin. The mean is given by the analytic model and the uncertainties are given by Eq. 5 and 6. Observational errors can also be added to this procedure, and we will consider them in detail in Sec. 3. Note that this process assumes that the dispersion along the stream is not substantially modified by the presence of the subhalo, which we will validate later in Sec. 2.3.
Figure 2: Dependence of some key observables, $\overline{\Delta\theta}$ and $\overline{\Delta v_{r}}$, on stream properties. The subhalo impact is the same as the right panel of Fig. 1. In each column, we vary one stream property and fix the others to demonstrate how a given property affects either the size of the error bars or the shape of the signal. The default stream properties are $\sigma_{\theta}=0.2$ deg, $r_{0}=10~\mathrm{kpc}$ and $\lambda=32~\mathrm{deg}^{-1}$. Signals here are only calculated from the analytic model and do not account for the uncertainties. Error bars here are only due to stream dispersion. When observational errors are included (Sec. 3), changing $r_{0}$ will also affect the error bars.
@@
and the dispersion in the radial, tangential and $z$ velocities are proportional to the circular velocity and the ratio of tidal radius to the orbital radius,
$\sigma_{v_{r}}\propto\sigma_{v_{t}}\propto\sigma_{v_{z}}\propto V_{c}\frac{r_{\mathrm{tidal}}}{r_{0}},$ (8)
where the tidal radius is (King, 1962)
@@
$\sigma_{r}\propto\left(\frac{M_{\mathrm{prog}}}{r_{0}}\right)^{1/3}r_{0},$ (10)
@@
From Eq. 5 and Eq. 6, the uncertainty also depends on the number of stars per bin $n$, or essentially the number of stars per unit angle $\lambda$. Following Fardal et al. (2015), the stream density in the central part of the stream is controlled by the total number of stream stars $N_{\mathrm{stars}}$ as,
@@
In addition to uncertainties, the detectability also depends on the signals (observables) of a subhalo impact. From Eqs. 1-3, we see that the observables depend on the stream’s distance to the Galactic Center $r_{0}$. The length of the stream is also important for the full signal to be present, since the impact is most relevant over a region of length $\sim w_{\mathrm{rel}}\sqrt{b^{2}+r_{s}^{2}}/w_{\perp}$. From Fardal et al. (2015), the angular length of the stream $l$ is controlled by the age of the stream $T_{\mathrm{form}}$ and the velocity dispersion in the tangential direction $\sigma_{v_{t}}$:
@@
$r_{0}$ distance to the Galactic Center 5405-40 kpc